OpenStudy (anonymous):

The ratio of the radii of two spheres is 3:1. The volume of the larger sphere is 729 in3. What is the volume of the smaller sphere? A. 81 in3 B. 54 in3 C. 243 in3 D. 27 in3

OpenStudy (loser66):

@terenzreignz

terenzreignz (terenzreignz):

formula for volume of a sphere is \[\Large V = \frac43\pi r^3\]

terenzreignz (terenzreignz):

If the radius of the smaller sphere is r, then the radius of the larger sphere should be...?

OpenStudy (loser66):

3r

terenzreignz (terenzreignz):

That's right... So the volume of the larger sphere should be \[\Large 729 = \frac43(3r)^3\]

OpenStudy (loser66):

where pi?

terenzreignz (terenzreignz):

Sorry, forgot. \[\Large 729 = \frac43\pi(3r)^3\]

terenzreignz (terenzreignz):

Now, we're looking for the volume of the smaller sphere, which is... \[\Large \frac{4}3\pi r^3\]

terenzreignz (terenzreignz):

What we can do here is... \[\Large 729 = \frac43\pi\cdot\color{red}{27\cdot r^3}\] Distribute the exponent 3, and remembering that \(3^3 = 27\)

OpenStudy (loser66):

got it, hehe. Thanks for the logic

terenzreignz (terenzreignz):

oh... good, you got it :D I'm so sleepy now, I'm sorry I went through it so quickly. Signing off. ------------------- Terence out :)

OpenStudy (loser66):

so funny that I am interest in other's post

terenzreignz (terenzreignz):

Indeed :)

OpenStudy (loser66):

sleep well, friend

OpenStudy (anonymous):

Thanks @Loser66 & @terenzreignz

OpenStudy (loser66):

yw